$SPAD/src/input multfact.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool multfact.output
)set message test on
)set message auto off
)clear all
--S 1 of 5
a := rootOf(a**2+a+1)
--R
--R
--R   (1)  a
--R                                                        Type: AlgebraicNumber
--E 1
--S 2 of 5
p := y*z**2 + a*z*x**2 + a*a*x*y**2
--R
--R
--R           2      2                2
--R   (2)  y z  + a x z + (- a - 1)x y
--R                                            Type: Polynomial(AlgebraicNumber)
--E 2
--S 3 of 5
factor(p,[a])
--R
--R
--R           2      2                2
--R   (3)  y z  + a x z + (- a - 1)x y
--R                                  Type: Factored(Polynomial(AlgebraicNumber))
--E 3
--S 4 of 5
b:=rootOf(b**2+1)
--R
--R
--R   (4)  b
--R                                                        Type: AlgebraicNumber
--E 4
--S 5 of 5

factor(x**2*y**2+u**2*v**2,[b])
--R
--R
--R   (5)  (x y - b u v)(x y + b u v)
--R                                  Type: Factored(Polynomial(AlgebraicNumber))
--E 5
)spool
)lisp (bye)

———-

References^{[1] nothing}

5